# Gaussian is All You Need: A Unified Framework for Solving Inverse Problems via Diffusion Posterior Sampling

### v1

@@ Abstract
Diffusion models can generate a variety of high-quality images by modeling complex data distributions. Trained diffusion models can also be very effective image priors for solving inverse problems. Most of the existing diffusion-based methods integrate data consistency steps within the diffusion reverse sampling process. The data consistency steps rely on an approximate likelihood function. In this paper, we show that the existing approximations are either insufficient or computationally inefficient. To address these issues, we propose a unified likelihood approximation method that incorporates a covariance correction term to enhance the performance and avoids propagating gradients through the diffusion model. The correction term, when integrated into the reverse diffusion sampling process, achieves better convergence towards the true data posterior for selected distributions and improves performance on real-world natural image datasets. Furthermore, we present an efficient way to factorize and invert the covariance matrix of the likelihood function for several inverse problems. We present comprehensive experiments to demonstrate the effectiveness of our method over several existing approaches.
@@ 1 Introduction
Diffusion-based models [16, 30, 33, 34] have recently gained attention due to their powerful generative ability by learning complex data distributions. A number of recent methods have used diffusion-based models as priors for solving inverse problems. [19, 41]. Diffusion Posterior Sampling (DPS) [9] and $\Pi$GDM [32] are notable examples that incorporate widely adopted posterior sampling schemes in the reverse diffusion process. Nevertheless, both methods are computationally inefficient as they require computing gradients through the diffusion model to compute the conditional score. Furthermore, DPS uses inexact likelihood covariance matrix estimates (even for simple Gaussian priors).
@@
In this paper, we argue that the Gaussian prior is all you need to achieve the best of both worlds, providing both covariance correction and computational efficiency. Our work begins by computing the distribution of measurements conditioned on intermediate outputs of the diffusion network for simple Gaussian priors. For a linear inverse problem with a Gaussian prior, the conditional distribution has an explicit Gaussian form. We show that this distribution can lead to the conditional distribution proposed in $\Pi$GDM although both methods make start from different assumptions. For isotropic covariance cases, our expression is equivalent to DPS up to a scalar constant (see Section 3.1). While the conditional distribution is generally intractable for non-Gaussian priors, the form obtained under the Gaussian assumption serves as an approximation that results in remarkable reconstruction performance. Another key advantage of this assumption is that it allows us to bypass gradient computation through the network when computing the score of the conditional distribution. Our experiments show that computing gradients through the network can be up to 2×2$\times$2 × costly.
@@
The Gaussian assumption for approximating the conditional score is already used, both implicitly and explicitly, in various diffusion-based posterior sampling methods. For instance, in $\Pi$GDM the conditional Gaussian assumption can easily be shown to be equivalent to the Gaussian prior assumption. As a consequence of this assumption, the method was able to introduce a covariance term in the likelihood. Another recent approach DDS [10] proposed a sampling scheme that does not compute gradients through the network. This work proposed to approximate the gradient computation using an orthogonal projection to the clean data manifold. However, to arrive at this simplification, DDS assumes that the prior has a Gaussian form with an infinite variance.
@@
In this paper, we introduce a Covariance Corrected Diffusion Posterior Sampling (CoDPS) method, which utilizes a likelihood approximation method with a covariance correction term that depends on the diffusion time-steps and the forward model. This approximation is exact for simple Gaussian priors, and we demonstrate that it can improve reconstruction performance in real-world restoration problems. Since the correction term is time-dependent, we can encode the uncertainty in our conditional guidance. Specifically, at the early stages of the diffusion process, the uncertainty in our estimates is high and gradually decreases as $t\to 0$. Our proposed method is time and memory efficient as it does not need to compute gradients through the network. Moreover since the correction term is high-dimensional, we employ efficient factorization and inversion during the sampling process to estimate the covariance. We consider a large class of inverse problems, including image inpainting, super-resolution, Gaussian and motion deblurring, and separable systems. We present a proof-of-concept experiment to illustrate that our method quickly converges to the true data covariance for a Gaussian mixture model. Through experimental evaluation on real-world natural image datasets, we demonstrate that our method enhances both reconstruction and perceptual quality. Additionally, our method does not compute gradients through the diffusion network and is more time- and memory-efficient.
@@ 2.1 Diffusion models
During training, we aim to learn $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$, which are used to predict $\bm{x}_{t-1}$ given $\bm{x}_{t}$ during the reverse sampling process. The training objective of a diffusion model minimizes the KL divergence between the joint distributions $q(\bm{x}_{0,\dots,T})$ and $p(\bm{x}_{0},\dots,T)$. The parameterization in [16] demonstrates that the trained denoiser $\bm{\epsilon}_{\theta}$ can be used to predict the mean $\bm{\mu}_{\theta}$ as follows
$\bm{\mu}_{\theta}(\bm{x}_{t},t)=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\bm{x}_{t})\right).$ (3)
Similarly, the covariance $\bm{\Sigma}_{\theta}(\bm{x}_{t},t)$ can be learned or set as a fixed parameter, such as $\sigma_{t}^{2}\bm{I}$. Some of the possible choices for $\sigma_{t}^{2}$ include $\beta_{t}$ or $\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ [16]. Once the mean predictor $\bm{\epsilon}_{\theta}$ and the covariance $\sigma_{t}^{2}\bm{I}$ are trained or determined, they can be used in the reverse sampling process.
Score SDE models [34] learn the score $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$ with a diffusion network $\bm{s}_{\theta}(\bm{x}_{t},t)$ using denoising score matching objectives [25, 26, 35]. Once the score network is trained, we can use it to obtain samples from regions of high probability. The connection between the denoiser network $\bm{\epsilon}_{\theta}(\bm{x}_{t})$ and the score function $\bm{s}_{\theta}(\bm{x}_{t},t)$ is given by $\bm{s}_{\theta}(\bm{x}_{t},t)={-\bm{\epsilon}_{\theta}(\bm{x}_{t})}/{\sqrt{1-\bar{\alpha}_{t}}}$ [34]. With this relationship, (3) can be considered as an unconditional estimation. In this process, we effectively move in the direction of $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$, which guides us towards regions of higher probability density in the data distribution. In Section 2.2, we show how this unconditional sample generation can be utilized to solve inverse problems, where we aim obtain samples from high-probability regions that are also consistent with our measurements.
@@ 2.2 Posterior sampling for inverse problems
where $\bm{y}$ represents measurements, $\mathcal{A}(\cdot)$ represents forward operator or a measurement matrix, and $\sigma_{n}$ represents noise. The maximum a posteriori (MAP) ${\bm{x}}_{\text{MAP}}$ estimate can be found by maximizing the sum of the likelihood $p(\bm{y}|\bm{x}_{0})$ and the data prior distribution $p(\bm{x}_{0})$. From equation (4), it is evident that $p(\bm{y}|\bm{x}_{0})$ is a Gaussian distribution. However, identifying a suitable prior $p(\bm{x}_{0})$ remains an active area of research. Pre-trained Diffusion models are suitable candidates as they learn the log prior density $\nabla_{\bm{x}}\log p(\bm{x})$. To utilize them for solving inverse problems, we can adapt the unconditional sampling in (3) for measurement-conditioned sampling. This involves incorporating a conditional score in the reverse sampling process as
@@
The main challenge then lies in modeling $p(\bm{x}_{0}|\bm{x}_{t})$. Using the reverse diffusion sampling process, we can approximate and sample from $p(\bm{x}_{0}|\bm{x}_{t})$. However, marginalizing over $\bm{x}_{0}$ is infeasible, as generating a single sample requires performing a complete reverse ancestral sampling. To circumvent this, DPS[9] proposed to approximation the conditional score $\bm{\kappa}_{t}$ as
@@
where $\hat{\bm{x}}_{0}(\bm{x}_{t})$ is given by (3) and $p(\bm{y}|\hat{\bm{x}}_{0})\simeq\mathcal{N}(\mathcal{A}(\hat{\bm{x}}_{0}),\sigma^{2}_{n}\bm{I})$. Alternatively, $\Pi$GDM [32] assumes $p(\bm{x}_{0}|\bm{x}_{t})\sim\mathcal{N}(\hat{\bm{x}}_{0},r_{t}^{2}\bm{I})$, where $r_{t}$ is a time- and data-dependent hyper-parameter, and approximates the conditional score as
@@
where $\bm{\Sigma}_{t}=r_{t}^{2}\bm{A}\bm{A}^{\mathsf{H}}+\sigma_{n}^{2}\bm{I}$ Both approximations in DPS and $\Pi$GDM have demonstrated remarkable results on various real-world datasets. However, the approximation in DPS may become inaccurate as $t$ approaches the total number of diffusion steps $T$. Additionally, both methods compute backpropagation through the diffusion network to evaluate $\partial\hat{\bm{x}}_{0}/\partial\bm{x}_{t}$, which can significantly increase the time and memory complexity of the sampling process. Another related work [24] introduced a covariance estimation method applicable to both $\Pi$GDM and DiffPIR-type conditional guidance methods. Denoising Diffusion Restoration Models (DDRMs) [19], on the other hand, use singular value decomposition (SVD) to perform diffusion in the spectral domain. However, the primary limitation of this method is it that relies on the SVD of the forward operator. DiffPIR [41] is another related method that integrates diffusion models in a plug-and-play manner. Other approaches, such as [8, 34, 7, 20, 21, 18], utilize pre-trained diffusion models. In this work, we demonstrate the connections between DPS, $\Pi$GDM, and DDS-type conditioning methods, and propose a unified diffusion posterior sampling algorithm. Our approach leverages implicit assumptions present in previous works and enhances computational efficiency by eliminating gradient computation through the network. While the focus our work is mainly unsupervised methods, diffusion models have also been used to solve several inverse problem in a fully-supervised manner [29, 27, 37, 38, 11].
@@ 3.1 Conditional score approximation
We revisit the conditional score approximation from DPS [9], stated as $p(\bm{y}|\bm{x}_{t})\approx\mathcal{N}(\bm{A}\hat{\bm{x}}_{0},\sigma_{n}^{2}\bm{I})$ for linear inverse problem under additive Gaussian measurement noise, where $\hat{\bm{x}}_{0}$ is the quantity given in (3). Intuitively, this approximation appears to be valid for small values of $t$ as the diffusion model learns the true data distribution $p(\bm{x}_{0})$ and the clean image estimates $\hat{\bm{x}}_{0}$ will accurately represent samples from this distribution. However, at the early stages of the reverse diffusion process $t\approx T$, the samples $\bm{x}_{t}$ will be close to an isotropic Gaussian distribution. The clean image estimates $\hat{\bm{x}}_{0}$ from $\bm{x}_{t}$ at these stages will not necessarily come from the the true prior distribution. This can be easily verified by looking at the estimates $\hat{\bm{x}}_{0}$ for different values of $t$, as shown in Figure 2. When $1000600<t<$ 1000, the estimates do not seem to resemble clean natural images (i.e samples from $p(\bm{x}_{0})$). Additionally, within the range $t\in[200,600]$, although the samples appear plausible, they are perceptually different from the ground truth image. More importantly, the uncertainly in the measurement $\bm{y}$ when conditioned on these estimates grows as $t$ increases and needs to be accounted for in the conditional variance. To see this more clearly, let us consider a simple prior Gaussian distribution given by $\bm{x}_{0}\sim\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$. Using the properties of marginal and conditional Gaussian distributions, we obtain
@@
As $t\to 0$, $\bar{\alpha}_{t}\to 1$ and $p(\bm{x}_{0}|\bm{x}_{t})\to\delta\left(\bm{x}_{0}-\frac{1}{\sqrt{\bar{\alpha}_{t}}}\bm{x}_{t}\right)$ (i.e., a delta distribution at the scaled noisy estimate $\bm{x}_{t}$). Similarity, as $t\to T$, where $T$ is the number of diffusion steps, $\bar{\alpha}_{t}\to 0$ and $p(\bm{x}_{0}|\bm{x}_{t})\to\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$, which is the prior data distribution. Note that the posterior mean provided by (8) matches the posterior mean obtained through Tweedie’s identity in (3) (see more details in Appendix A.1).
@@
The forward diffusion in (1) ensures that $p(\bm{x}_{t}|\bm{x_{0}})$ is a normal distribution. We note that for a given $\bm{x}_{t}$, $r_{t}$ in (7) is a constant. Theorem 1 further shows that the joint distribution $p(\bm{x}_{0},\bm{x}_{t})$ is a bivariate normal distribution. From this we can easily deduce that the marginals $p(\bm{x}_{0})$ and $p(\bm{x}_{t})$ are also Gaussian. Following the arguments above, we can conclude that the Gaussian assumption on $p(\bm{x}_{0}|\bm{x}_{t})$ implies that $p(\bm{x}_{0})$ itself is a Gaussian.

### v2

@@ Abstract
Diffusion models can generate a variety of high-quality images by modeling complex data distributions. Trained diffusion models can also be very effective image priors for solving inverse problems. Most of the existing diffusion-based methods integrate data consistency steps by approximating the likelihood function within the diffusion reverse sampling process. In this paper, we show that the existing approximations are either insufficient or computationally inefficient. To address these issues, we propose a unified likelihood approximation method that incorporates a covariance correction term to enhance the performance and avoids propagating gradients through the diffusion model. The correction term, when integrated into the reverse diffusion sampling process, achieves better convergence towards the true data posterior for selected distributions and improves performance on real-world natural image datasets. Furthermore, we present an efficient way to factorize and invert the covariance matrix of the likelihood function for several inverse problems. Our comprehensive experiments demonstrate the effectiveness of our method over several existing approaches. Code available at https://github.com/CSIPlab/CoDPS.
@@ 1 Introduction
Diffusion-based models [17, 34, 38, 39] have recently gained attention due to their powerful generative ability by learning complex data distributions. A number of recent methods have used diffusion-based models as priors for solving inverse problems [21, 48]. Diffusion Posterior Sampling (DPS) [9] and $\Pi$GDM [37] are notable examples that incorporate widely adopted posterior sampling schemes in the reverse diffusion process. Nevertheless, both methods are computationally inefficient as they require computing gradients through the diffusion model to compute the conditional score. Furthermore, DPS uses inexact likelihood covariance matrix estimates (even for simple Gaussian priors).
@@
In this paper, we show that the Gaussian prior assumption is all that is needed to achieve the best of both worlds, offering both covariance correction and computational efficiency. Building on this assumption, we propose a novel and simple sampling framework that is not only computationally efficient but also achieves performance that is competitive with, and in some cases surpasses, existing methods. Our work begins by computing the distribution of measurements conditioned on intermediate outputs of the diffusion network for simple Gaussian priors. For a linear inverse problem with a Gaussian prior, the conditional distribution has an explicit Gaussian form. This distribution has the same form as the conditional distribution proposed in $\Pi$GDM, even though both methods start from different assumptions (see Sec. 3.1). Furthermore, for the isotropic covariance case, our conditional distribution is equivalent to the one used in DPS, up to a scalar constant (see Sec. 3.1). While the conditional distribution is generally intractable for non-Gaussian priors, the form obtained under the Gaussian assumption serves as an approximation that results in remarkable reconstruction performance. Figure 1 presents some examples that illustrate how our method successfully recovers fine details that other methods fail to restore. Another key advantage of this assumption is that it allows us to bypass gradient computation through the network when computing the score of the conditional distribution. Our experiments show that bypassing gradients through the network can provide at least 2×2$\times$2 × reduction in the computational cost.
@@
The Gaussian assumption for approximating the conditional score is already used, both implicitly and explicitly, in different diffusion-based posterior sampling methods. For instance, in $\Pi$GDM the conditional Gaussian assumption can be shown to be equivalent to the Gaussian prior assumption. As a consequence of this assumption, the method was able to introduce a covariance term in the likelihood. A recent method, DDS [10], introduces a novel sampling strategy that entirely bypasses gradient computations through the network. Instead, it approximates the gradient by projecting orthogonally onto the clean data manifold. This approach significantly simplifies the sampling process by assuming that the prior distribution is Gaussian with infinite variance.
@@
In this paper, we introduce a Covariance Corrected Diffusion Posterior Sampling (CoDPS) method, which utilizes a likelihood approximation method with a covariance correction term that depends on the diffusion time steps and the forward model. This approximation is exact for Gaussian priors, and we demonstrate that it can improve reconstruction performance in real-world restoration problems. Since the correction term is time-dependent, we can encode the uncertainty in our conditional guidance. Specifically, at the early stages of the diffusion process, the uncertainty in our estimates is high and gradually decreases as $t\to 0$.
We summarize our main contributions as follows.
  • We first demonstrate that by making a simple choice of a Gaussian prior, we get a conditional score that is accurate under this prior and eliminates the need for expensive gradient computations through the diffusion network.
  • Building on the Gaussian score approximation, we propose Covariance Corrected Diffusion Posterior Sampling (CoDPS), a novel framework for solving inverse problems. This framework significantly reduces time and memory requirements.
  • We further accelerate CoDPS by efficient factorization and inversion of the high-dimensional covariance matrices for a family of inverse problems.
  • We validate the accuracy of CoDPS through a proof-of-concept experiment on a Gaussian mixture and demonstrate its effectiveness on real-world image datasets. Our experiments show that CoDPS achieves competitive performance with remarkable computational efficiency.
@@ 2.1 Diffusion models
During training, we aim to learn $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$, which are used to predict $\bm{x}_{t-1}$ given $\bm{x}_{t}$ during the reverse sampling process. The training objective of a diffusion model minimizes the KL divergence between the joint distributions $q(\bm{x}_{0},\dots,\bm{x}_{T})$ and $p(\bm{x}_{0},\dots,\bm{x}_{T})$. The parameterization in [17] demonstrates that the trained denoiser $\bm{\epsilon}_{\theta}$ can be used to predict the mean, which we denote as $\hat{\bm{x}}_{0}$:
${\hat{\bm{x}}_{0}}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\bm{x}_{t})\right).$ (3)
Similarly, the covariance $\bm{\Sigma}_{\theta}(\bm{x}_{t},t)$ can be learned or set as a fixed parameter, such as $\sigma_{t}^{2}\bm{I}$. Once the mean predictor $\bm{\epsilon}_{\theta}$ and the covariance $\sigma_{t}^{2}\bm{I}$ are trained or determined, they can be used in the reverse sampling process.
Score SDE models [39] learn the score $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$ with a diffusion network $\bm{s}_{\theta}(\bm{x}_{t},t)$ using denoising score matching objectives [29, 30, 40]. Once the score network is trained, we can use it to obtain samples from regions of high probability. The score function can be approximated by a denoising model that learns the underlying probability distribution of the training data, as shown in previous work [18, 39]. The relationship between the denoiser network $\bm{\epsilon}_{\theta}(\bm{x}_{t})$ and the score function $\bm{s}_{\theta}(\bm{x}_{t},t)$ is expressed as
$\bm{s}_{\theta}(\bm{x}_{t},t)=\frac{-\bm{\epsilon}_{\theta}(\bm{x}_{t})}{\sqrt{1-\bar{\alpha}_{t}}}.$ (4)
This equation highlights how the output of the denoiser is linked to the gradient of the log likelihood of the observed data. By establishing this connection, we can view (3) as an unconditional estimation informed by the prior implicit in the denoiser. In Section 2.2, we show how this unconditional sample generation can be utilized to solve inverse problems, where we aim to obtain samples from high-probability regions that are also consistent with our measurements.
@@ 2.2 Posterior sampling for inverse problems
where $\bm{y}$ represents measurements, $\mathcal{A}(\cdot)$ represents forward operator or a measurement matrix, and $\sigma_{n}$ represents noise. The maximum a posteriori (MAP) ${\bm{x}}_{\text{MAP}}$ estimate can be found by maximizing the sum of the likelihood $p(\bm{y}|\bm{x}_{0})$ and the data prior distribution $p(\bm{x}_{0})$. From equation (5), it is evident that $p(\bm{y}|\bm{x}_{0})$ is a Gaussian distribution. However, identifying a suitable prior $p(\bm{x}_{0})$ remains an active area of research. Pre-trained Diffusion models are suitable candidates as they learn the log prior density $\nabla_{\bm{x}}\log p(\bm{x})$. To utilize them for solving inverse problems, we build upon the unconditional denoising prediction in (3) to develop a measurement-conditioned sampling procedure. This involves incorporating a conditional score in the reverse sampling process as
@@
The term $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$ in (6) can be approximated using a pre-trained diffusion model $\bm{s}_{\theta}(\bm{x}_{t},t)$. However, as noted in [37, 9], $p(\bm{y}|\bm{x}_{t})$ is intractable. This issue is discussed in DPS [9] and $\Pi$GDM [37], where it is highlighted that there is no analytical form available for $p(\bm{y}|\bm{x}_{t})$. To understand why this is the case, let us consider
@@
where $\hat{\bm{x}}_{0}$ is given by (3) and $p(\bm{y}|\hat{\bm{x}}_{0})\simeq\mathcal{N}(\mathcal{A}(\hat{\bm{x}}_{0}),\sigma^{2}_{n}\bm{I})$. Alternatively, $\Pi$GDM [37] assumes $p(\bm{x}_{0}|\bm{x}_{t})\sim\mathcal{N}(\hat{\bm{x}}_{0},r_{t}^{2}\bm{I})$, where $r_{t}$ is a time- and data-dependent hyper-parameter, and approximates the conditional score as
@@
where $\bm{\Sigma}_{t}=r_{t}^{2}\bm{A}\bm{A}^{\mathsf{H}}+\sigma_{n}^{2}\bm{I}$ Both approximations in DPS and $\Pi$GDM have demonstrated remarkable results on various real-world datasets. However, the approximation in DPS may become inaccurate as $t$ approaches the total number of diffusion steps $T$. Additionally, both methods compute backpropagation through the diffusion network to evaluate $\partial\hat{\bm{x}}_{0}/\partial\bm{x}_{t}$, which can significantly increase the time and memory complexity of the sampling process.
Several recent works have proposed strategies to eliminate the expensive gradient computation through the diffusion network during inverse problem solving. One such approach is DMPS [24], which introduces a closed-form approximation of the conditional likelihood by assuming an uninformative prior on $p(\bm{x}_{0})$. This assumption allows the method to avoid backpropagation entirely. Notably, our Gaussian prior also represents DMPS as a special case when $\sigma_{0}^{2}\to\infty$. DDNM [41] introduces a zero-shot framework for solving linear inverse problems by leveraging the structure of the forward operator. It decomposes the solution into two orthogonal components: the row space and the null space of the operator $\bm{A}$. The row-space component is obtained directly via the pseudoinverse projection $\bm{A}^{\dagger}\bm{y}$, ensuring strict data consistency. The null-space component is refined by iteratively denoising with a pretrained diffusion model and projecting the result onto the null space using $(\mathbf{I}-\bm{A}^{\dagger}\bm{A})\hat{\bm{x}}_{0}$. However, this method is unstable in the presence of measurement noise. To address this, the authors proposed DDNM+, which includes a range-space correction term. In our experiments, we observed poor performance of DDNM+ on deblurring tasks in the presence of noise, despite the correction term. ReSample [35] is another method that leverages pretrained latent diffusion models (LDMs). It avoids gradient computation by enforcing hard data consistency in the latent space through constrained optimization, a process referred to as a ReSample timestep. During the early stages of the reverse sampling process, the method performs unconditional DDIM updates without applying any ReSample steps. Resampling occurs every $k$ iterations, where $k$ is typically set to 10. Each ReSample step involves two stages: first, minimizing the reconstruction error $\|\bm{y}-\bm{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$, where $\bm{z}$ is the latent variable initialized from $\bm{\hat{z}}_{0}$ (the latent-space equivalent of the predicted mean in Eq. (3)) and $\mathcal{D}$ is the decoder that maps latent vectors back to the image space. Once the minimizer is obtained, the second stage performs stochastic resampling to generate the new latent variable $\bm{z}_{t}$. Despite only applying sampling at certain steps, ReSample has significantly higher runtime. Additionally, it requires both an encoder and a decoder to map data to and from the latent space, which adds complexity and limits its applicability in settings without pretrained auto-encoder models. Other representative method is DDRM [21], which performs spectral-domain posterior sampling but relies on SVD of the forward operator and assumes linearity. DiffPIR [48] follows a plug-and-play strategy, applying pretrained diffusion models as proximal denoisers without explicitly modeling the posterior. Recent extensions [28] propose covariance estimation modules to enhance conditioning accuracy in such frameworks.
A complementary line of work, referred to as variational inference-based methods, approximates the posterior distribution with a simpler, tractable distribution instead of seeking for a closed-form expression for the conditional score  [11]. RED-Diff [23] proposes to minimize the KL divergence between the true conditional posterior $p(\bm{x}_{0}\mid\bm{y})$ and a variational distribution $q(\bm{x}_{0}|\bm{y})=\mathcal{N}(\bm{\mu},\sigma^{2}\bm{I})$. The mean $\bm{\mu}$ is iteratively updated by minimizing the KL objective, while the variance of the variational distribution is assumed to be small $\sigma^{2}\approx 0$. While RED-Diff and CoDPS share the goal of modeling the posterior that utilize simple Gaussian forms, the two approaches differ significantly in formulation and implementation. RED-Diff treats posterior sampling as a stochastic optimization problem and relies on iterative updates guided by the denoiser, whereas CoDPS derives a closed-form conditional score correction under a Gaussian prior.
Other approaches, such as [8, 39, 7, 22, 25, 20], utilize pre-trained diffusion models. In this work, we demonstrate the connections between DPS, $\Pi$GDM, and DDS-type conditioning methods, and propose a unified diffusion posterior sampling algorithm. Our approach leverages implicit assumptions present in previous works and enhances computational efficiency by eliminating gradient computation through the network. While the focus of our work is mainly on unsupervised diffusion based methods, several deep-learning based inverse problem in a fully-supervised manner [33, 31, 43, 44, 12, 45].
@@ 3.1 Conditional score approximation
We revisit the conditional score approximation from DPS [9], stated as $p(\bm{y}|\bm{x}_{t})\approx\mathcal{N}(\bm{A}\hat{\bm{x}}_{0},\sigma_{n}^{2}\bm{I})$ for linear inverse problem under additive Gaussian measurement noise, where $\hat{\bm{x}}_{0}$ is the quantity given in (3). Intuitively, this approximation appears to be valid for small values of $t$ as the diffusion model learns the true data distribution $p(\bm{x}_{0})$ and the clean image estimates $\hat{\bm{x}}_{0}$ will accurately represent samples from this distribution. However, at the early stages of the reverse diffusion process $t\approx T$, the samples $\bm{x}_{t}$ will be close to an isotropic Gaussian distribution. The clean image estimates $\hat{\bm{x}}_{0}$ from $\bm{x}_{t}$ at these stages will not necessarily come from the true prior distribution. This can be easily verified by looking at the estimates $\hat{\bm{x}}_{0}$ for different values of $t$, as shown in Figure 2. When $1000600<t<$ 1000, the estimates do not seem to resemble clean natural images (i.e samples from $p(\bm{x}_{0})$). Additionally, within the range $t\in[200,600]$, although the samples appear plausible, they are perceptually different from the ground truth image. More importantly, the uncertainty in the measurement $\bm{y}$ when conditioned on these estimates grows as $t$ increases and needs to be accounted for in the conditional variance. To see this more clearly, let us consider a simple prior Gaussian distribution given by $\bm{x}_{0}\sim\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$. Using the properties of marginal and conditional Gaussian distributions, we obtain
@@
As $t\to 0$, $\bar{\alpha}_{t}\to 1$ and $p(\bm{x}_{0}|\bm{x}_{t})\to\delta\left(\bm{x}_{0}-\frac{1}{\sqrt{\bar{\alpha}_{t}}}\bm{x}_{t}\right)$ (i.e., a delta distribution at the scaled noisy estimate $\bm{x}_{t}$). Similarly, as $t\to T$, where $T$ is the number of diffusion steps, $\bar{\alpha}_{t}\to 0$ and $p(\bm{x}_{0}|\bm{x}_{t})\to\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$, which is the prior data distribution. Note that the posterior mean provided by (10) matches the posterior mean obtained through Tweedie’s identity in (3) (see more details in Appendix A.1).
@@
We provide the proof of Theorem 1 in Appendix A.2. The forward diffusion in (1) ensures that $p(\bm{x}_{t}|\bm{x_{0}})$ is a normal distribution. We note that for a given $\bm{x}_{t}$, $r_{t}$ in (9) is a constant. Theorem 1 further shows that the joint distribution $p(\bm{x}_{0},\bm{x}_{t})$ is a bivariate normal distribution. From this we can easily deduce that the marginals $p(\bm{x}_{0})$ and $p(\bm{x}_{t})$ are also Gaussian. Following the arguments above, we can conclude that the Gaussian assumption on $p(\bm{x}_{0}|\bm{x}_{t})$ implies that $p(\bm{x}_{0})$ itself is a Gaussian.